# Almost commuting self-adjoint operators and measurements

### v3

@@ Abstract
We study the problem when an -tuple of self-adjoint operators in an infinite dimensional separable Hilbert space with small commutators is close to an -tuple of commuting self-adjoint operators on We give an affirmative answer to the problem when the synthetic-spectrum and the essential synthetic-spectrum are close. In the case that we show that a pair of almost commuting self-adjoint operators in an infinite dimensional separable Hilbert space is close to a commuting pair of self-adjoint operators if and only if a corresponding Fredholm index vanishes outside of an essential synthetic-spectrum. Examples are also exhibited that, in general, the answer to the problem when is negative even the associated Fredholm index vanishes. This is an attempt to solve a problem proposed by David Mumford related to quantum theory and measurements.
@@ 1 Introduction
In the First International Congress of Basic Science held in 2023, David Mumford gave the opening plenary lecture on Consciousness, robots and DNA ([33]). At the very beginning, he cautioned the audience that there would be no mathematics in his talk. Nevertheless, by the very end of this magnificent lecture, when he ended his talk on DNA being a measuring instrument opening a Pandora’s box, he proposed a problem in -algebras about non-commuting self-adjoint operators to be approximated by commuting ones.
A version of this problem proposed by Mumford may be reformulated below: Let and be a positive integer. Is there a constant such that the following statement holds? If is any separable Hilbert space and are self-adjoint operators on (with ) such that
@@
In the case that and is any finite dimensional Hilbert space (no bound on the dimension) the exactly the same problem is known as von-Neumann-Kadison-Halmos problem for almost commuting self-adjoint matrices (see [11] and [12]). In that case, the problem has an affirmative solution [20] (see also [8] and [13], and, for a fixed dimension, see [32]). In an infinite dimensional Hilbert space, as observed by Mumford in his new book (p.182), the answer in general is negative (see also Example 4.6 of [8]) in general. The affirmative solution for almost commuting self-adjoint matrices led the further study in the subject of weakly semi-projectivity in -algebras. It also has a deep impact on the Elliott program of classification of amenable -algebras (see [7], [24], [26] and [28]). The Mumford’s problem opens a new door to the future applications of this part of -algebra theory in previous unexpected areas such as DNA and artificial intelligence ([34]).
@@
For a pair of self-adjoint operators, we also have a more definite answer for some Hilbert modules.
@@
The second condition in (e 1.7) means that a Fredholm index vanishes (see Definition 2.4). Note that both case (ii) and case (iii) are generalization of case (i). In case (ii), is stable (whenever is not a finitely generated projective Hilbert module), and is simple (like the Calkin algebra). In case (ii), we keep the feature that is simple and has real rank zero.
Case (i) in Theorem 1.3 is a more definite result than that of Theorem 1.1 in the case We actually prove a slightly more general result (see Theorem 5.3 below). One should also compare (i) in Theorem 1.3 to Theorem 1.1 of [17] (see Remark 5.6). In the case that however, the normal elements may not be approximated by normal elements in with finite spectrum. We also show that condition of the vanishing index associated with the pair of operators is necessary. Since it is well-known that spectrum (and the essential spectrum) of a bounded operator is notoriously unstable regarding perturbation, it is necessary to consider some version of -spectrum (which is usually larger than the spectrum). For we introduce the notion of -synthetic-spectrum for -tuples of self-adjoint elements in a unital -algebra (see Definition 2.7). We will also show that, unfortunately, the version of Mumford’s problem formulated above has a negative solution in general even the associated Fredholm index vanishes.
In quantum mechanics, one often heard “when the commutators go to zero, we recover the classical system”. If “recover” means the observables are approximated by commuting observables, then example in (iii) of Chapter 14 of [34] (Example 4.6 of [8]), Proposition 5.5 and Proposition 6.4 all show that, mathematically, there are serious (topological) obstacles prevent this type of statements from being true. Nevertheless Theorem 1.1 with some interpretation in Remark 6.6 may provide an alternative statement.
@@ 2 Preliminaries
Let be a -algebra and be a Hilbert (right) module over (see [15]), or Hilbert -module. Denote by the -algebra of all bounded module maps on with adjoints. A rank one module map is a bounded module map of the form for all (and fixed ), where is the -value inner product on Denote by the linear span of rank one module maps and the (norm) closure of Following Theorem 1 of [15], we identify with
@@
Fix an integer Define
has only finitely many points.
@@
To see this, let There is with such that There is such that Write and If then Therefore Hence Therefore
@@
for any which has value on an open ball with the center (for some ) and the radius Then we say that is an -near-spectrum of the -tuple We write
If, moreover, is a unital homomorphism, then we say is an - spectrum of the -tuple
Let be an -near-spectrum and be a -near -spectrum for respectively. Suppose that then, by the definition, is also an -near-spectrum. In particular, if then
@@
We often use the -near spectrum as it associated with a c.p.c. map, whenever the -tuple sufficiently almost commutative as shown in the next proposition (2.11). But it is not unique and depends on the choice of the c.p.c. map. On the other hand, the -synthetic-spectrum is always defined and, in the computation world, it is computable. Their relationship is described in Proposition 2.15 below (see also Proposition 2.16). In practice, it is quite possible that it is more difficult to compute the spectrum of an operator where than to compute the norm of
@@
(1) and
@@
Suppose that, for case (1), (or for case (2), ). Then we obtain an and a sequence of unital -algebras and a sequence of -tuples with (), such that
@@
To show (1), consider Fix Choose such that then Hence, since is a monomorphism,
@@
Recall that Hence there exists such that
@@
It follows that for all In other words, has -synthetic-spectrum containing This proves (1).
For case (2), we obtain, by the Choi-Effros Theorem ([4]), a c.p.c. map such that Write where each is a c.p.c map, Then
@@
Let and Choose such that (the second part of) Lemma 2.12 holds for above and (in place of ). Choose such that, for any unital c.p.c. map (any unital -algebra ), that
@@
By the choice and applying (the second part of) Lemma 2.12, we have
@@
It follows that (keeping in mind of (e 2.45))
@@
Combining (e 2.49) with (e 2.47), we obtain
@@
We conclude that, by (e 2.50) and (e 2.46),
| (e 2.51) | |||
@@
Let be a metric space (we only consider the case that ). Recall that the Hausdorff distance of two compact subsets of is defined by
| (e 2.54) |
Let be the set of all non-empty compact subsets of Then is a compact metric space with the metric
@@
(1) Suppose that is a unital -algebra and with () such that has a non-empty -near-spectrum If is also a non-empty -near-spectrum of and then
| (e 2.55) |
@@
Note that is a generating set of Let be given by Lemma 2.12 (for and ) and be given by Corollary 2.13. Choose
Now suppose that and are -near-spectra of the -tuple Let be c.p.c maps such that
| (e 2.57) | |||
@@
| (e 2.59) |
for any with for all whenever (if ) or (if ). Let We may assume that Then By (e 2.59),
| (e 2.60) |
By (e 2.58) and the choice of applying Lemma 2.12,
@@
Since if we have Therefore
@@
| (e 2.67) |
For (2), let us keep notation above. Then is immediate, and that and that follow from part (1). It remains to show that
@@
| (e 2.72) | |||
| (e 2.73) | |||
| (e 2.74) |
for any with if for some Choose We may assume that Hence It follows from (e 2.74) that
| (e 2.75) |
On the other hand, by the choice of (e 2.58), (e 2.73), and applying Lemma 2.12, we have
| (e 2.76) |
Thus
@@
Hence It follows that
| (e 2.78) |
Hence ∎
@@
| (e 2.80) |
and a sequence of numbers Put and Note that and Since is compact, without loss of generality, we may assume that for some compact subset in as and
Let and be the quotient map. Put Then, by (e 2.80), is normal. Let Then defined by for is a unital monomorphism. By the Choi-Effros Lifting Theorem ([4]), there is a sequence of unital c.p.c. maps such that Let be as in Proposition 2.15. Then, since is a homomorphism, for all large is a -near-spectrum for Applying Proposition 2.15, we may assume that, for all
| (e 2.83) |
Note that, by (e 2.82) and (e 2.83), Let Note that Hence there are such that ()
@@ 3 Purely infinite simple -algebras
An even - brick with corner is a subset of of the form
| (e 3.86) |
For the remaining of this paper, all -bricks are even and has the corner An open -brick is the interior of the -brick By boundary of the -brick we mean
A -brick combination is a finite union of -bricks. The -brick combination has the following easy property: If are two distinct -bricks of then
We think Proposition 3.2 are well known.
@@
If is convex, is contractive. Hence and
@@
| (e 3.91) |
Since we have established that and are all finitely generated, we conclude that is also finitely generated. This ends the induction.
@@
Since has property (F), we may write where each is path connected. Let be such that and if Put (). By considering for each we reduce the general case to the case that has only one path connected component. Put for some Then we have a splitting short exact sequence
@@
| (e 3.95) |
Define by for all Then one computes that, by (e 3.93),
| (e 3.96) |
@@
| (e 3.103) |
Define for all Then is a unital homomorphism from to with finite dimensional range and, by (e 3.103), (e 3.100) and (e 3.98), for all
Finally, since in as we arranged earlier, by the definition of and the “Moreover” part of the statement also holds. ∎
@@
Let be a compact subset of Suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism. If is any compact subset with property (F) such that where then there is a unital injective homomorphism such that
| (e 3.105) |
@@
| (e 3.106) |
Suppose that is a compact subset of such that Since is compact, there is with such that
| (e 3.107) |
Put and Then There is a finite -net such that where Then
Choose a non-zero projection with Since is a purely infinite simple -algebra, there is a unital embedding Applying Theorem 2.8 of [16], one obtains a unital embedding Note in There is a partial isometry such that and Define by for all Then is a unital injective homomorphism.
Choose mutually orthogonal non-zero projections such that Define by Define by Then is injective. One computes that
| (e 3.108) |
@@
| (e 3.110) |
We may write for all where and are mutually orthogonal non-zero projections with By choosing a better approximation if necessary, without loss of generality, we may assume that is -dense in Recall that is -dense in and -dense in For each choose exactly one such that Then, we may assume where is a set of distinct points such that
| (e 3.111) |
@@
| (e 3.115) |
Then, combining with (e 3.110) and (e 3.112), for all
| (e 3.116) | |||||
@@
| (e 3.117) |
Then is a unital injective homomorphism. Moreover, for all by (e 3.109), the definition of above and by (e 3.116),
| (e 3.118) | |||||
To see the second formula in (e 3.105), we may write, from (e 3.117), since factors through
| (e 3.119) |
@@
Suppose that is a compact subset of and suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism. If is any compact subset with property (F) such that then there is a unital injective homomorphism such that
| (e 3.120) |
@@
Let be a compact metric space which is a a compact subset of (for integer ) and be a unital purely infinite simple -algebra. Suppose that is a unital injective homomorphism. Then, for any and any finite subset there exists such that, for any integer with there exists a unital injective homomorphism such that
| (e 3.122) |
where is the quotient map by restriction, is a finite subset such that (and where is the -brick cover of ).
@@
Put Choose any with and By Theorem 3.6, we obtain a -brick combination and a unital injective homomorphism such that
| (e 3.126) |

### v4

@@ Abstract
We study the problem when an -tuple of self-adjoint operators in an infinite dimensional separable Hilbert space with small commutators is close to an -tuple of commuting self-adjoint operators on We give an affirmative answer to the problem when the synthetic-spectrum and the essential synthetic-spectrum are close. Examples are also exhibited that, in general, the answer to the problem when is negative even the associated Fredholm index vanishes. This is an attempt to solve a problem proposed by David Mumford related to quantum theory and measurements.
@@ 1 Introduction
At the First International Congress of Basic Science held in 2023, David Mumford delivered the opening plenary lecture on Consciousness, Robots and DNA ([33]). He began by cautioning the audience that there would be no mathematics in his talk. Nevertheless, by the very end of this magnificent lecture—which culminated with the idea of DNA as a measuring instrument opening a Pandora’s box— he proposed an unexpected problem in -algebras: determining when almost-commuting self-adjoint operators can be approximated by commuting ones.
A version of this problem proposed by Mumford may be reformulated as follows: Let and be a positive integer. When is there a constant such that the following statement holds? For any separable Hilbert space and self-adjoint operators (with ) satisfying
@@
In the case that and is any finite dimensional Hilbert space (no bound on the dimension) the exactly the same problem is known as von-Neumann-Kadison-Halmos problem for almost commuting self-adjoint matrices (see [11] and [12]). In this setting, the problem has an affirmative solution [20] (see also [8] and [13], and [32] for fixed dimensions). However, in an infinite dimensional Hilbert space as Mumford noted in his recent book (p.182) – and as was previously known – the answer is negative in general (see also Example 4.6 of [8]).
The affirmative solution for almost commuting self-adjoint matrices has spurred further research in the study of weak semi-projectivity in -algebras. It also has a deep impact on the Elliott program of classification of amenable -algebras (see [7], [24], [26] and [28]). Moreover, Mumford’s problem opens new avenues for applying this branch of -algebra theory to previously unexpected fields, such as DNA research and artificial intelligence ([34]).
@@
While we are mostly interested in the case that we also looks the special case when For a pair of self-adjoint operators, we have a more definite answer for some Hilbert modules.
@@
The second condition in (e 1.7) means that a Fredholm index vanishes (see Definition 2.4). Note that both cases (ii) and (iii) generalize case (i). In case (ii), is stable (whenever is not a finitely generated projective Hilbert module), and is simple (analogous to the Calkin algebra). Moreover, in case (ii), retains its simiplicity while has real rank zero.
Case (i) in Theorem 1.3 is a more definite result than that of Theorem 1.1 for In fact, we prove a slightly more general result (see Theorem 5.3 below). One should also compare (i) in Theorem 1.3 to Theorem 1.1 of [17] (see Remark 5.6). Theorem 1.1 of [17] provides a more definitive estimate of the distance from an element to the set of normal elements with finite spectrum in -algebras with real rank zero. However, when normal elements may not be approximated by normal elements in with finite spectrum. Thus, cases (ii) and (iii) in Theorem 1.3 cover scenarios where almost normal elements are close to normal elements that cannot be approximated normal elements with finite spectra — complementing the results in [17].
We further demonstrate that the vanishing index condition for the operator pair is necessary. Given well-known instablity of the spectrum (and essential spectrum) of bounded operators under perturbations, it becomes essential to consider some version of -spectrum (typically larger than the spectrum). For we introduce the concept of -synthetic-spectrum for -tuples of self-adjoint elements in a unital -algebra (see Definition 2.7). We will also show that, unfortunately, the version of Mumford’s problem formulated above has a negative solution in general— even the associated Fredholm index vanishes when
In quantum mechanics, one often encouters the statement, “when the commutators go to zero, we recover the classical system”. If “recover” means the observables are approximated by commuting observables, then example in (iii) of Chapter 14 of [34] (Example 4.6 of [8]), Proposition 5.5 and Proposition 6.4 all reveal significnat topological obstructions to such claims. Nevertheless Theorem 1.1, interpreted as in Remark 6.6, may offer an alternative perpective.
@@ 2 Preliminaries
Let be a -algebra and a Hilbert (right) module over (or Hilbert -module; see [15]). Denote by the -algebra of all bounded module maps on with adjoints. A rank-one module map is a bounded module map of the form for all (and fixed ), where is the -valued inner product on Denote by the linear span of rank-one module maps and the (norm) closure of Following Theorem 1 of [15], we identify with
@@
Fix Fix an integer Define
has only finitely many points. Most of the time, we choose
@@
To see this, let There is with such that There is such that Write and If then
Therefore Hence Therefore
@@
for any which has value on an open ball with the center (for some ) and the radius Then we say that is an -near-spectrum of the -tuple We may write
for convenince. The reader is recommended to read Remark 2.10 for clarification.
If is, additionally, a unital homomorphism, then we say is an - spectrum of the -tuple
@@
Here are some clarifications.
1. It should be noted that is not uniquely determined and depends on the choice of the c.p.c. map. Writing simply means that there exists a non-empty compact subset and a unital c.p.c. map satisfying conditions (i), (ii) and (iii) of Definition 2.9. This convention is adopted for notational convenience. Moreover, if we write we mean that no such and exist which satisfy (i), (ii) and (iii).
2. However, we will later show that -near spectra are unique up to an - neighborhood when the -tuple selfadjoint operators are almost commuting within as elaborated in Proposition 2.15.
3. On the other hand, the -synthetic-spectrum is always uniquely defined and—importantly– computable in the context of computations. Their relationship is described in Proposition 2.15 below (see also Proposition 2.16).
4. Let be an -near-spectrum and be a -near -spectrum for respectively. Suppose that then, by the definition, is also an -near-spectrum. In particular, if then
5. In practice, computing the spectrum of an operator (where ) may be more challenging than computing the norm of
@@
(1) and,
(2) i.e., there exists a non-empty compact subset and a c.p.c. map satisfies condition (i), (ii) and (iii) in Definition 2.9.
@@
Suppose that the lemma is false. Then we obtain an and a sequence of unital -algebras and a sequence of -tuples with (), such that
@@
To show (1), consider (see Definition 2.7). Fix Choose such that then Hence, since is a monomorphism,
@@
Recall that Hence there is an infinite subset such that, for all
@@
It follows that for all In other words, has -synthetic-spectrum containing for all This contradicts “ ”. Thus (1) follows.
For (2), we obtain, by the Choi-Effros Theorem ([4]), a c.p.c. map such that Write where each is a c.p.c map, Then
@@
Let and For choose such that (the second part of) Lemma 2.12 holds for above and (in place of ). Choose such that, for any unital c.p.c. map (any unital -algebra ), that
@@
Define by for all By the choice and applying (the second part of) Lemma 2.12 to we obtain
@@
In other words,
@@
Note that It follows that (keeping in mind of (e 2.45))
@@
Combining (e 2.50) with (e 2.47), we obtain
| (e 2.51) |
We conclude that, by (e 2.51) and (e 2.46),
@@
Let be a metric space (we only consider the case that is compact such as the case that ). Recall that the Hausdorff distance of two compact subsets of is defined by
| (e 2.55) |
Let be the set of all non-empty compact subsets of Then is a compact metric space with the metric when is compact.
@@
(1) Suppose that is a unital -algebra and with () such that has a non-empty -near-spectrum If is also a non-empty -near-spectrum of then and
| (e 2.56) |
@@
Note that is a generating set of Let be given by Lemma 2.12 (for and as well as ) and be given by Corollary 2.13. Choose
Now suppose that and are -near-spectra of the -tuple Let be unital c.p.c maps such that
@@
| (e 2.59) | |||
| (e 2.60) | |||
@@
for any with for all whenever (for ), or (for ).
Let We may assume that Then By (e 2.61),
@@
Define () by and for all By (e 2.59) and applying Lemma 2.12, we have, for
| (e 2.63) |
Then, by (e 2.62), we obtain, for
| (e 2.64) |
Since if we must have Therefore
| (e 2.65) |
@@
| (e 2.70) |
Since we conclude that The same argument shows that
Next, let By (e 2.67),
| (e 2.71) |
Hence Thus
| (e 2.72) |
It follows that
| (e 2.73) |
Hence
| (e 2.74) |
This completes the proof of part (1).
For (2), let us keep notation above. Then is immediate.
Suppose that Then that follow from part (1). It remains to show that
@@
| (e 2.75) | |||
| (e 2.76) | |||
@@
for any with if for some
@@
and a sequence of numbers Put and Note that and Since is compact, without loss of generality, by passing to a subsequence, we may assume that for some compact subset in as and
Let and be the quotient map. Put Then, by (e 2.84), is normal. Let Then defined by for is a unital monomorphism. By the Choi-Effros Lifting Theorem ([4]), there is a sequence of unital c.p.c. maps such that Let be as in Proposition 2.15. Then, since is a homomorphism, for all large is a (non-empty) -near-spectrum for Applying Proposition 2.15, we may assume that, for all
| (e 2.87) |
Note that, by (e 2.86) and (e 2.87), Let Note that Hence there are such that ()
| (e 2.88) |
@@ 3 Purely infinite simple -algebras
An even - brick with corner (with – see Definition 2.6) is a subset of of the form
| (e 3.90) |
For the remaining of this paper, all -bricks are assumed to be even and have their corner An open -brick is the interior of the -brick The boundary of a -brick is defined as
A -brick combination is a finite union of -bricks. Since two distinct - bricks have distinct corners, a -brick combination has the following easy property: If are two distinct -bricks of then
We believe that Proposition 3.2 to be well known.
@@
If is convex, then is contractive. Hence and
@@
| (e 3.95) |
Since we have established that and are all finitely generated abelian groups, we conclude that is also finitely generated. This ends the induction.
@@
Since has property (F), we may write where each is path connected and is finitely generated (), Let be such that and if Put (). By considering for each we reduce the general case to the case that has only one path connected component. Note that, if originally, is not connected, then we do not need to consider the “Moreover” part of the theorem. If is connected, then we need to point out that
So now we assume that is connected. Put for some Then we have a splitting short exact sequence
@@
| (e 3.99) |
Define by for all Then one computes that, by (e 3.97),
| (e 3.100) |
@@
| (e 3.107) |
Define, for all
| (e 3.108) |
Then is a unital homomorphism from to with finite dimensional range and, by (e 3.107), (e 3.104) and (e 3.102), for all
Finally, since in as we arranged earlier, by (e 3.108), the “Moreover” part of the statement also holds. ∎
@@
Let be a compact subset of Suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism. If is any compact subset with property (F) such that then there is a unital injective homomorphism such that
| (e 3.110) |
@@
| (e 3.111) |
Suppose that is a compact subset of such that which has property (F). Since is compact, there is with such that
| (e 3.112) |
Put and Then There is a finite -net such that where Then Note that we assume that if
Choose a non-zero projection with in and Since is a purely infinite simple -algebra, there is a unital embedding Applying Theorem 2.8 of [16], one obtains a unital embedding Note in There is a partial isometry such that and Define by for all Then is a unital injective homomorphism.
Choose mutually orthogonal non-zero projections such that Define by Then in Define by for all Then is injective. One computes that
| (e 3.113) |
@@
| (e 3.115) |
We may write for all where and are mutually orthogonal non-zero projections in with By choosing a better approximation if necessary, without loss of generality, we may assume that is -dense in Recall that is -dense in and -dense in For each choose exactly one such that Then, we may assume where is a set of distinct points such that
| (e 3.116) |
@@
| (e 3.120) |
Then, combining with (e 3.115) and (e 3.117), for all
| (e 3.121) | |||||
@@
| (e 3.122) |
Then is a unital injective homomorphism. Moreover, for all by (e 3.114), the definition of and above and by (e 3.121),
| (e 3.123) | |||||
To see the second formula in (e 3.110), we may write, from (e 3.122), since factors through
| (e 3.124) |
@@
Suppose that is a compact subset of is a unital purely infinite simple -algebra and is a unital injective homomorphism. If is any compact subset with property (F) such that then there is a unital injective homomorphism such that
| (e 3.125) |
@@
Let Then, for any and any finite subset there exists satisfying the following: Suppose that is a purely infinite simple -algebra and is a unital injective homomorphism, where is a compact subset. For any integer with there exists a unital injective homomorphism such that
| (e 3.127) |
where is the quotient map by restriction, is a finite subset such that (and where is a -brick cover of ).
@@
Put Choose any with and By Corollary 3.8, we obtain a -brick combination and a unital injective homomorphism such that
| (e 3.131) |